(4x+5)/4x = 5 ?????????????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = (4x+5)/4x = 4x/4x + 5/4x = 1 +5/4x ≠5 = RHS
The correct statement is (4x+5)/4x = 1 + (5/4x)

Thanks